a rectangular pane of glass used as a window measures 200 mm by 950 mm and its 7mm thick .calculate the heat loss through the pane if the outside tempeture is 6t celcius and the inside temperature is 25 celcius


1
Expert's answer
2021-02-11T08:54:11-0500

Thermal conductivity of glass is k=0.84W/(m·ºC).

The heat loss thru the glass is


"q=\frac{kA\Delta T}{d}=\frac{0.84·0.2\cdot0.95(25-6)}{0.007}=433.2\text{ W}."